# Occam’s LGS: A simple approach for Language Gaussian Splatting

### v1

@@ 3.3 Filtering Noisy Gaussians
After the forward rendering pass, we identify and filter out Gaussians that have no contributions across all views. This filtering effectively reduces noise while reducing the number of Gaussians significantly. Through this process, we obtain a refined set of semantic Gaussians $\mathcal{G}^{s}=\{{G_{i}^{s}}\}_{i=1}^{M}$, where $M\ll N$, representing only the Gaussians that meaningfully contribute to the scene’s semantic representation.
@@
7:     $\alpha_{i}\leftarrow 0$Accumulated weights}
8:  end for
9:  Feature Uplifting Phase:
10:  for each camera view $s\in\mathcal{S}$ do
11:     for each Gaussian $G_{i}$ in view frustum do
12:        $\alpha_{i}^{s},\mathbf{p}_{i}^{s}\leftarrow\text{Rasterize}(G_{i}|s)$Project and rasterize}
13:        $\mathbf{f}_{i}^{s}\leftarrow$ 2D feature map for view $s$
14:        $\mathbf{f}_{i}\leftarrow\mathbf{f}_{i}+\alpha_{i}^{s}\mathbf{f}_{i}^{s}$Sampled feature}
15:        $\alpha_{i}\leftarrow\alpha_{i}+\alpha_{i}^{s}$Alpha blending value}
16:     end for
17:  end for
18:  for each Gaussian $G_{i}\in\mathcal{G}$ do
19:     $\mathbf{f}_{i}\leftarrow\mathbf{f}_{i}/\alpha_{i}$Feature Normalization}
20:  end for
21:  return  $\mathcal{G}$
@@ 4 Experiments
We evaluate our method against other state-of-the-art 3D Language Gaussian Splatting methods on open vocabulary segmentation and localization tasks, as well as compare their computational efficiency metrics.
Figure 3: PCA visualization of semantic features on LERF Dataset. (a) Original RGB rendering. (b) PCA visualization of our rendered semantic feature maps. (c) Initial semantic feature maps extracted using CLIP from SAM segmentations. Note that colors are not aligned across different pairs of images due to independent PCA projections.
@@ 4.1 Dataset
To assess the effectiveness of our method, we evaluate it on two widely used datasets: LERF [13] and 3D-OVS [17]. The LERF dataset features everyday indoor scenes with diverse objects and we perform quantitative evaluations following the on the commonly used scenes Ramen, Figurines, Teatime, and Waldo Kitchen according to the LERF evaluation protocol [13]. For these scenes, we evaluate both mean Intersection over Union (mIoU) and localization accuracy. The 3D-OVS dataset contains long-tail objects in diverse backgrounds, and we evaluate the mIoU on five scenes: sofa, lawn, bed, room, and bench. For evaluation, we utilize the annotations provided by [24] for both datasets.
@@ 4.2 Implementation Details
Our implementation of Occam’s LGS is built on gsplat [38], a CUDA-accelerated Gaussian rasterization library that supports high-dimensional feature rendering. For each scene, we train a 3D Gaussian Splatt for 30,000 iterations using the common Gaussian Splatting [12] parameters, resulting in an average number of 2,500,000 Gaussians for RGB representation across scenes. To ensure a fair comparison, we follow the 2D language feature map generation method from [24]. For language feature extraction and segmentation, our pipeline integrates OpenCLIP ViT-B/16 [26] and SAM ViT-H [14]. To capture semantic information hierarchically, we generate 2D feature maps at three levels: whole, part, and subpart, where each level represents distinct semantic granularity. Following the feature extraction, we use our method to uplift the 2D language features to 3D scenes using training views only, and evaluate the results on separate testing views, withheld during training of the semantic representation. All experiments and runtime measurements are conducted on a single NVIDIA A6000 GPU.
@@ 4.3 Open Vocabulary Semantic Segmentation
Given a 3D scene and a language query describing an object or region of interest, the task of open vocabulary semantic segmentation is to identify and generate masks for all relevant regions that match a query description. We evaluate our approach against existing methods on two datasets: LERF and 3D-OVS, with quantitative results shown in Tables 2 and 3 respectively and qualitative results shown in Figure 3 and 4. On the LERF dataset, under identical experimental settings as LangSplat, our approach demonstrates a substantial improvement of 9.9% mIoU. Notably, our method shows advantages, especially in complex scenes like Waldo Kitchen, where the camera captures an entire room rather than focusing on a specific area surrounded by camera views. Thus, these challenging scenarios come with sparse view coverage, compared to the size of the scene.
@@
Figure 4: Qualitative comparison of relevancy score visualization. Trained with the same feature maps generated using CLIP and SAM.
@@ 4.4 Localization
We further evaluate on the language-guided localization task, where given a language query, the goal is to localize the described object in 3D space. Following [13], we select the pixel with the highest relevancy score as the location of the object. Benchmarking against existing methods, our approach achieves comparable localization accuracy, as shown by Table 4.
@@ 4.5 Computational Efficiency
We evaluate the computational efficiency of our method against existing approaches in Figure 1. Similar to ours, LangSplat and GOI both build on pre-trained 3D Gaussian models, and we trained GOI using identical feature maps as LangSplat at whole object, part, and subpart granularities. For these existing methods, we report their semantic feature training times. LangSplat takes 67 minutes to train three levels of feature maps, while GOI takes 60 minutes of training time. With our approach, language features can be integrated in 16.35 seconds without setup (including checkpoint loading and saving), and 31.93 seconds with setup. In contrast, 3DVLM cannot utilize pre-trained Gaussians and thus, the reported time reflects their complete joint training of both reconstruction and semantics, which takes 65 minutes [22]. Thus, our method, which allows to use pre-trained visual Gaussians achieves semantic feature integration two orders of magnitude faster than other methods, while delivering superior mIoU performance across test datasets.
To analyze the scalability of our method, we evaluate the relationship between training time and the number of Gaussians present in a scene. For this purpose, we test our approach on four scenes, each containing a different number of Gaussians due to varying scene complexity and resolutions. The results depicted in Figure 1 demonstrate that our method scales well and maintains high computational efficiency even with large scenes containing up to 4 Million Gaussians, indicating good scalability in practice.
@@ 4.6 Ablation Study
We validate the key components of our method with an ablation studies depicted in Table 5 and Figure 5. First, we ablate our fundamental choice to model the forward rendering process with the weighted feature aggregation strategy. Removing it, while keeping the Gaussian filtering process (w/o wtd. agg), shows that weighted aggregation improves the fusion of multi-view features due to the accurate modeling. Second, we analyze the impact of filtering redundant Gaussians that are not activated during rendering. Removing filtering (w/o filter) decreases performance, which supports the choice to filter inactive Gaussians, besides the observation that it enhances the smoothness and improves computational efficiency. Finally, we compare to a baseline not using weighting and filtering and thus, perform basic back-projection. This reduces performance by $4.2\%$, and altogether demonstrates that both components contribute meaningfully to our method’s performance.
@@ 4.7 Application: Object Insertion
Figure 6: Visualization of Object Insertion: (a) Object Selection: A specific object is selected in the Figurines scene (b) Object Insertion: The selected object is transferred and integrated into the Teatime scene (c) Feature Analysis: PCA visualization of the feature map to show the object’s integration in the semantic space (d) Query Visualization: Heatmap showing the response to the query “pink icecream".
A major advantage of the proposed method is the deep integration of dense 3D language features in 3D Gaussian Splatting and the inherent simplicity of its feature representation. This allows us to directly profit from any advances in 2D language feature representations as well as from the open-set generalization capability of these methods. Therefore, our approach has the potential to enable a range of downstream 3D tasks, especially related to scene editing and interaction.
We demonstrate this capability on object insertion, where an object is taken from an existing 3D Gaussian Splatt and integrated in another scene. While the current state-of-the-art LangSplat would require finetuning of the scene-specific language autoencoder, as well as the retraining of the whole feature embedding, our representation can be augmented with the new object similarly to any other Gaussian splat. As illustrated in Figure 6, we select and extract an ice cream cone (represented by its Gaussians) from the Figurines scene of LERF [13]. By simply copying the object’s Gaussians together with their parameters and semantic features, the new object seamlessly integrates into the Teatime scene while preserving its semantic features.
To verify the successful integration, we query the modified scene with "pink icecream". The resulting relevancy heatmap in Figure 6) (d) confirms that our method maintains the semantic understanding of the inserted object in its new context. This demonstrates that, unlike existing methods that require scene-specific dimensional reduction techniques, our approach enables direct manipulation of semantic features, making it more flexible for practical scene editing operations while preserving semantic querying capabilities.
@@ 5 Discussion and Conclusion
In this work, we presented an optimization-free approach for augmenting Gaussian Splatting with language features. The proposed method shows state-of-the-art results with a single iteration over all views, and avoids rendering and loss computation in the feature space.
@@
  • High-dimensional language features can directly be used in Gaussian splatting by utilizing a training free optimization approach.
@@
Based on the experiments and results presented in this work, we can answer the following questions about our work and it’s relation to existing works using feature compression:
Is feature compression or spatial mapping necessary for Language Gaussian Splatting? No. Avoiding the explicit rendering of language features and loss computation in the 2D space together with a training-free approach makes the direct use of high-dimensional language features feasible. Avoiding compressions algorithms significantly reduces training time and helps generalization of Language Gaussian Splatting.
Should we stop using low-dimensional Language Gaussian Splatting? No. Low dimensional features are justified in certain scenarios with consideration of their constraints. Though our method enables the direct use of uncompressed features, low-dimensional features can reduce the memory footprint of any 3D representation that relies on language features and speed up rendering of 3D featuremaps. However, it needs to be considered that using low-dimensional embeddings contradicts generalization across scenes, if compression approaches are trained scene-specifically.
Are there applications where low-dimensional language features are beneficial? Yes. For scenarios where the domain of scenes is constrained, no scene editing is required, and storage and rendering cost exceeds the cost of training a feature compressor, low dimensional features are advantageous. Thus, the optimal feature resolution is strongly application dependent and scales especially with the domain of potential scenarios.
Overall, it is important to acknowledge that our approach is complementary to existing works in Language Gaussian Splatting and enables a new direction to train the 3D language representation. Due to this complementary nature, Occam’s LGS can be combined with existing feature compression approaches, with the potential to speed up rendering and reduce storage cost. The right combination of approaches depends on the application and expected usecase of the 3D representation, which is a direction we hope future work will explore.
@@
6 Evaluation

### v2

@@ 3.4 Filtering Noisy Gaussians
After the forward rendering pass, we identify and filter out Gaussians that have no contributions across all views. This filtering effectively reduces noise while reducing the number of Gaussians significantly. Besides, this statistical outlier elimination is consistent with our probabilistic framework, as Gaussians with minimal visibility have high uncertainty in their feature estimates. Through this process, we obtain a refined set of semantic Gaussians $\mathcal{G}^{s}=\{{G_{i}^{s}}\}_{i=1}^{M}$, where $M\ll N$, representing only the Gaussians that meaningfully contribute to the scene’s semantic representation.
@@
7:     $w_{i}\leftarrow 0$Accumulated weights}
8:  Feature Uplifting Phase:
9:  for each camera view $s\in\mathcal{S}$ do
10:     for each Gaussian $G_{i}$ in view frustum do
11:        $w_{i}^{s},\mathbf{p}_{i}^{s}\leftarrow\text{Render}(G_{i}|s)$Project and Render}
12:        $\mathbf{f}_{i}^{s}\leftarrow$ 2D feature map for view $s$
13:        $\mathbf{f}_{i}\leftarrow\mathbf{f}_{i}+w_{i}^{s}\mathbf{f}_{i}^{s}$Sampled feature}
14:        $w_{i}\leftarrow w_{i}+w_{i}^{s}$Alpha blending value}
15:  for each Gaussian $G_{i}\in\mathcal{G}$ do
16:     $\mathbf{f}_{i}\leftarrow\mathbf{f}_{i}/w_{i}$Feature Normalization}
17:  return  $\mathcal{G}$
@@ 4 Experiments
We evaluate our method against other state-of-the-art 3D Language Gaussian Splatting methods on open vocabulary segmentation and localization tasks, as well as compare their computational efficiency metrics. Additionally, we analyze how our method scales across different feature dimensions, examining the trade-offs between dimensionality, computational requirements, and segmentation quality.
Figure 3: PCA visualization of semantic features on LERF Dataset. (a) Original RGB rendering. (b) PCA visualization of our rendered semantic feature maps. (c) Initial semantic feature maps extracted using CLIP from SAM segmentations.
@@ 4.1 Dataset
To assess the effectiveness of our method, we evaluate it on two widely used datasets: LERF [13] and 3D-OVS [17]. The LERF dataset features everyday indoor scenes with diverse objects and we perform quantitative evaluations on the commonly used scenes Ramen, Figurines, Teatime, and Waldo Kitchen according to the LERF evaluation protocol [13]. For these scenes, we evaluate both open-vocabulary segmentation (mIoU) and localization accuracy. The 3D-OVS dataset contains long-tail objects in diverse backgrounds, and we evaluate segmentation performance on five scenes: sofa, lawn, bed, room, and bench. For evaluation, we utilize the annotations provided by [23] for both datasets.
@@ 4.2 Implementation Details
Our implementation of Occam’s LGS is built on gsplat [37], a CUDA-accelerated Gaussian rendering library that supports high-dimensional feature rendering. For each scene, we train a 3D Gaussian Splatt for 30,000 iterations using the common Gaussian Splatting [12] parameters, resulting in an average number of 2,500,000 Gaussians for RGB representation across scenes. To ensure a fair comparison, we follow the 2D language feature map generation method from [23]. For language feature extraction and segmentation, our pipeline integrates OpenCLIP ViT-B/16 [25] and SAM ViT-H [14]. To capture semantic information hierarchically, we generate 2D feature maps at three levels: whole, part, and subpart, where each level represents distinct semantic granularity. Following the feature extraction, we use our method to uplift the 2D language features to 3D scenes using training views only, and evaluate the results on separate testing views, withheld during training of the semantic representation. All experiments and runtime measurements are conducted on a single NVIDIA A6000 GPU.
@@ 4.3 Open Vocabulary Semantic Segmentation
Given a 3D scene and a language query describing an object or region of interest, the task of open vocabulary semantic segmentation is to identify and generate masks for all relevant regions that match a query description. We evaluate our approach against existing methods on two datasets: LERF and 3D-OVS, with quantitative results shown in Tables 2 and 3 respectively and qualitative results shown in Figure 3 and 4. On the LERF dataset, under identical experimental settings as LangSplat, our approach demonstrates a substantial improvement of 9.9% mIoU. Notably, our method shows advantages, especially in complex scenes with sparse view coverage. An example of this is Waldo Kitchen, where the camera captures an entire room rather than focusing on a specific area that is fully surrounded by camera views.
@@
Figure 4: Qualitative comparison of relevancy score visualization.
@@ 4.4 Localization
We further evaluate on the language-guided localization task, where, given a language query, the pixel with the highest relevancy score is selected, following [13]. Benchmarking against existing methods, our approach achieves competitive localization accuracy, as shown by Table 4.
Adding feature compression as in existing methods, increases localization accuracy of our approach to 84.1. This is due to the metric only measuring the closest matching pixel, while ignoring the feature distribution and object shape, making it prone to noise. We provide detailed experiments in this setting in the supplementary material.
@@
4.5 Robustness to Feature Dimension Variations
In the following, we demonstrate the advantages of different feature dimensions that Occam’s LGS enables and assess the respective scalability and performance. For these experiments, we 1) used an autoencoder to compress the CLIP features, 2) uplifted them to 3D using our method, 3) rendered the compressed features, and 4) decompressed features to 512D using the autoencoder. Table 5 shows the IoU across different feature dimensions on LERF datasets. Notably, segmentation performance improves to 64.5 %mIoU with 64D features, outperforming the LangSplat [23] baseline by $13.1\%$ mIoU. This confirms that feature compression filters noise by limiting the feature space to objects present in the 3d scene, at the cost of limited scene editing capability. Since our method supports arbitrary feature dimensions, selecting this operating point is straightforward and can be used depending on the task at hand.
Dimensions Ramen Figurines Teatime Waldo Kitchen Overall
16-D 48.4 57.7 75.9 59.9 60.5
32-D 53.6 61.8 78.6 62.6 64.1
64-D 51.7 63.4 79.0 63.7 64.5
128-D 52.1 62.4 74.3 62.6 62.9
256-D 53.0 62.6 74.3 61.1 62.7
512-D 51.0 58.6 70.2 65.3 61.3
Table 5: Average mIoU with varying feature dimensions on LERF scenes.
@@ 4.6 Computational Efficiency
We evaluate the computational efficiency of our method against existing approaches in Figure 1 measuring their semantic Gaussian training times. Similar to ours, LangSplat and GOI both allow the integration of pre-trained 3D Gaussian models, while 3DVLM requires joint uplifting and reconstruction. LangSplat, GOI, and 3DVLM take 67, 60, and 65 minutes to uplift three levels of features respectively [21]. With our approach, language features can be uplifted in 16.4 seconds without setup, and 31.9 seconds with setup (checkpoint loading and saving). Overall, our method, achieves feature uplifting two orders of magnitude faster than existing methods, while delivering superior performance across test datasets and avoiding feature compression.
We assess our method’s scalability through a thorough analysis of runtime, storage requirements, rendering speed, and GPU memory to uplift one feature level across four scenes of varying complexity. Additionally, we evaluate scaling across different feature dimensions, providing insights into the range of practical deployment scenarios.
The results depicted in Figure 5(d) demonstrate that our method scales well and maintains high computational efficiency even with large scenes containing up to 4 Million Gaussians, indicating its practical scalability across varying scene complexities. The memory requirement, shown by Figure 5(a) stays below 14GB for 512D CLIP features, which is met by most modern GPUs. A key advantage of our method is its flexibility to support features of any dimension, enabling the selection of the optimal feature size for specific applications. From Figure 5(c), we observe that rendering speed decreases as feature dimension increases, presenting a clear trade-off between feature expressiveness and computational performance. Storage requirements similarly scale with feature dimensionality, shown by Figure 5(b). These findings highlight the adaptability of our approach, enabling users to balance quality, memory usage, and rendering speed based on their available resources and application requirements.
Figure 5: Comprehensive analysis of computational efficiency metrics across different feature dimensions, showing rendering speed (FPS), storage size, GPU memory usage, and runtime performance as Gaussian size and frame numbers vary.
@@ 4.7 Ablation Study
We validate the key components of our method with an ablation study depicted in Table 6 and Figure 6. First, we ablate our fundamental choice to model the forward rendering process with the weighted feature aggregation strategy. Removing it, while keeping the Gaussian filtering process (w/o wtd. agg), shows that weighted aggregation improves the fusion of multi-view features due to accurate modeling.
Second, we analyze the impact of filtering redundant Gaussians that are not activated during rendering. Removing filtering (w/o filter) decreases performance, which supports the choice to filter inactive Gaussians. This is because our method uplifts 2D features to a Gaussian only if its center projects onto the corresponding pixel, making it efficient but sparse for occluded Gaussians. Since we threshold occluded Gaussians with negligible influence $w_{i}^{s}<10^{-4}$ during training and inference, visual noise can occur. This happens because occluded Gaussians with minimal central influence may still affect neighboring pixels during rendering, but without receiving proper feature updates due to our sparse projection approach. This feature update-rendering influence mismatch creates semantic inconsistencies, so filtering occluded Gaussians improves both consistency and computational efficiency.
Finally, we compare to a baseline not using weighting and filtering(w/o wtd. agg + filter) and thus, perform basic back-projection. This reduces performance by $4.2\%$ and altogether demonstrates that both components contribute meaningfully to our method’s performance.
@@ 4.8 Application: Object Insertion
Figure 7: Visualization of Object Insertion: (a) Object Selection: An object is selected in the Figurines scene (b) Object Insertion: The object is transferred and integrated into the Teatime scene (c) Feature Analysis: PCA visualization of the feature map to show the object’s integration in the semantic space (d) Query Visualization: Heatmap showing the response to the query “pink icecream”.
Our method’s key advantage is integrating dense 3D language features into Gaussian Splatting with a simple representation. This allows us to profit from their strong open-set generalization capability and the advances in 2D language feature representations. This has the potential to enable a range of downstream 3D tasks, especially related to scene editing and interaction.
We demonstrate this capability on object insertion, where an object is taken from an existing 3D Gaussian Splat and integrated into another scene. While the current state-of-the-art LangSplat would require finetuning of the scene-specific language autoencoder, as well as the retraining of the whole feature embedding, our representation can be augmented with the new object similarly to any other Gaussian splat by simple copy-paste editing. As illustrated in Figure 7, we select and extract an ice cream cone (represented by its Gaussians) from the Figurines scene of LERF [13]. By simply copying the object’s Gaussians together with their parameters and semantic features, the new object seamlessly integrates into the Teatime scene while preserving its semantic features. Then, we query the modified scene with ”pink ice-cream”. The resulting relevancy heatmap in Figure 7) (d) confirms that our method maintains the semantic understanding of the inserted object in its new context. This demonstrates that our approach enables direct semantic feature manipulation without scene-specific dimensional reduction, offering flexibility for practical editing while preserving semantic querying capabilities.
@@ 5 Discussion and Conclusion
In this work, we presented a theoretically grounded and optimization-free approach for efficient Language Gaussian Splatting. The proposed method shows state-of-the-art results with a single iteration over all views and avoids rendering and loss computation in the feature space.
@@
  • High-dimensional language features can directly be used in Gaussian splatting by utilizing a training-free optimization approach.
@@
Based on the experiments and results presented in this work, we can answer the following questions about our work and its relation to existing works using feature compression:
Is feature compression or spatial mapping necessary for Language Gaussian Splatting? No. Avoiding the explicit rendering of language features and loss computation in the 2D space together with a training-free approach makes the direct use of high-dimensional language features feasible.
Should we stop using low-dimensional Language Gaussian Splatting? No. Low-dimensional features are justified in certain scenarios despite constraints. Though our method enables the direct use of uncompressed features, low-dimensional alternatives can reduce memory usage, and accelerate 2D feature map rendering, and filter noise. However, it needs to be considered that scene-specific compression approaches may contradict generalization across scenes.
Are there applications where low-dimensional language features are beneficial? Yes. For scenarios where the domain of scenes is constrained, no scene editing is required, and storage and rendering cost exceeds the cost of training a feature compressor, low dimensional features are advantageous. Thus, the optimal feature resolution is strongly application-dependent and scales, especially in the domain of potential scenarios.
Overall, it is important to acknowledge that our approach is complementary to existing works in Language Gaussian Splatting and enables a new direction to train the 3D language representation. Due to this complementary nature, Occam’s LGS can be combined with existing feature compression approaches, with the potential to speed up rendering and reduce storage costs. The right combination of approaches depends on the application and expected use case of the 3D representation, which is a direction we hope future work will explore.
@@
6 3D Feature Uplifting as Maximum Likelihood Estimation
Our weighted feature uplifting formula is strongly grounded in the forward rendering process from a probabilistic perspective: Given the task of estimating 3d features $\mathbf{f}\in\mathbb{R}^{d}$ from a set of observed 2D features $\mathbf{F}^{2D}\in\mathbb{R}^{d}$, the Maximum-a-Posteriori (MAP) estimator of 3D features $\mathbf{f}$ is
$\arg\max_{\mathbf{f}}p(\mathbf{f}|\mathbf{F}^{2D})=\arg\max_{\mathbf{f}}p(\mathbf{F}^{2D}|\mathbf{f})\frac{p(\mathbf{f})}{p(\mathbf{F}^{2D})},$ (8)
which forms the basis of our method.
Preliminaries
Given the LGS setting, we make the following key assumptions about the feature uplifting problem.
  1. (i)
    Uniform Prior: The language features in CLIP space have uniform distribution, which is motivated by the efficient use of the CLIP feature space, due to training of a large amount of diverse data
    $p(\mathbf{f})\propto\text{constant}.$ (9)
  2. (ii)
    Conditional Independence of pixels: Pixel observations are conditionally independent given the surface representation and 3D features $\mathbf{f}$
    $p(\mathbf{F}^{2D}|\mathbf{f})=\prod_{v=1}^{V}\prod_{p\in P_{v}}p(\mathbf{f}_{p,v}^{2D}|\mathbf{f}).$ (10)
    where $\mathbf{f}_{p,v}^{2D}$ denotes the 2D feature at pixel $p$ in view $v$.
  3. (iii)
    Gaussian feature noise: Viewpoint dependency of CLIP features can be modeled by a Gaussian distribution of feature noise
    $\epsilon_{p,v}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I}).$ (11)
  4. (iv)
    Alpha Blending Weights: Alpha blending weights applied to each Gaussian $i$ at pixel $p$ in view $v$ can be interpreted as measures of occlusion between separate Gaussians in a given camera view
    $w_{i,p,v}=\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}).$ (12)
Occam’s LGS as a MAP estimator
Our approach formulates feature uplifting as a MAP estimation problem, as stated in Equation 8. With the use of CLIP as language features, a model trained of large amounts of data is chosen. Under this condition, we make the assumption of a uniformly distributed feature space in Equation 9. Under this condition, the LGS estimator is equivalent to the Maximum Likelihood (ML)
$\arg\max_{\mathbf{f}}p(\mathbf{F}^{2D}|\mathbf{f}).$ (13)
This estimator can be evaluated separately for each 2D feature $\mathbf{f}_{p,v}^{2D}$ at pixel $p$ in view $v$ under conditional independence from Equation 10, given the surface representation and features $\mathbf{f}$ from the pretrained 3D Guassians
$\arg\max_{\mathbf{f}}p(\mathbf{F}^{2D}|\mathbf{f})=\arg\max_{\mathbf{f}}\prod_{v=1}^{V}\prod_{p\in P_{v}}p(\mathbf{f}_{p,v}^{2D}|\mathbf{f})$ (14)
where $P_{v}$ is the set of pixels in view $v$ on which Gaussian influences.
A major challenge in all LGS approaches is the viewpoint-dependent noise of 2D language features due to varying object appearance. Our approach formulates the ML estimation of LGS from noisy 2D features as IID normal distributed $\epsilon_{p,v}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I})$, representing the inherent uncertainty in the 2D features.
Hence, the 2D feature at pixel $p$ in view $v$, obtained from rendering the 3D Guassian features, is
$\mathbf{f}_{p,v}^{2D}=\sum_{k=1}^{R}w_{k,p,v}\mathbf{f}_{k}+\epsilon_{p,v},$ (15)
where $R$ is the number of Gaussians along the ray, $\epsilon_{p,v}$ is the language feature noise, and $w_{k,p,v}$ is the alpha blending weight for Gaussian $k$ at pixel $p$ in view $v$ defined in Equation 12
To approach the ML estimation problem $\arg\max_{\mathbf{f}}p(\mathbf{F}^{2D}|\mathbf{f})$, we transform the problem to the log-likelihood domain
$\log p(\mathbf{F}^{2D}|\mathbf{f})=\log\prod_{v=1}^{V}\prod_{p\in P_{v}}p(\mathbf{f}_{p,v}^{2D}|\mathbf{f})=$ (16)
$-\frac{1}{2\sigma^{2}}\sum_{v,p}\left\|\mathbf{f}_{p,v}^{2D}-\sum_{k=1}^{R}w_{k,p,v}\mathbf{f}_{k}\right\|^{2}+\text{const},$
and solve for the minimum by taking the derivative w.r.t. the 3D language feature $\mathbf{f}_{i}$ of Gaussian $i$
$\frac{\partial\log p}{\partial\mathbf{f}_{i}}=\frac{1}{\sigma^{2}}\sum_{v,p}w_{i,p,v}\left(\mathbf{f}_{p,v}^{2D}-\sum_{k=1}^{R}w_{k,p,v}\mathbf{f}_{k}\right)=0.$ (17)
This leads to a formula where the weighted sum of observed 2D features (left side) relates to the interaction of Gaussians during the rendering process (right side)
$\Rightarrow\sum_{v,p}w_{i,p,v}\mathbf{f}_{p,v}^{2D}=\sum_{k=1}^{R}\sum_{v,p}w_{i,p,v}w_{k,p,v}\mathbf{f}_{k}.$ (18)
Practical Simplification
Since our approach utilizes pretrained Gaussian Splats, we can make the assumption of sparse influence; i.e. for most views and pixels, one dominant Gaussian has significant influence (weight close to 1), which means that the rendering is dominated by non-overlapping Gaussians, leading to the following:
  1. (a)
    Negligible Cross-Talk: The product of weights from different Gaussians at the same pixel can be neglected compared to the dominant Gaussians, given the whole set of views $w_{i,p,v}w_{k,p,v}\approx 0$ for $i\neq k$. This eliminates the cross-Guassian terms in our equation.
  2. (b)
    Linearization: Since most weights are approximately binary (either close to 0 or close to 1), $w_{i,p,v}^{2}$ can be linearized as $\sum w_{i,p,v}^{2}\approx\sum w_{i,p,v}$, given it is used inside a sum.
With these properties of Gaussian Splats, the system of equations simplifies and can be written as the closed-form solution used in Ocamm’s LGS
$\boxed{\mathbf{f}_{i}=\frac{\sum\limits_{v=1}^{V}\sum\limits_{p\in P_{v}}w_{i,p,v}\mathbf{f}_{p,v}^{2D}}{\sum\limits_{v=1}^{V}\sum\limits_{p\in P_{v}}w_{i,p,v}}.}$ (19)
This formula has an intuitive interpretation: the optimal 3D feature for each Gaussian is a weighted average of the 2D features observed at pixels where the Gaussian has influence, with weights proportional to the Gaussian’s contribution to each pixel.
7 Robustness to Feature Dimension Variation
We also evaluate the localization accuracy across different feature dimensions on LERF dataset using the same feature compression method detailed in Sec 4.5. As shown by Table 7, localization accuracy improves to 84.1 when feature dimension reduces to 32D, indicating that feature compression effectively reduces noisy signals.
Dimensions Ramen Figurines Teatime Waldo Kitchen Overall
16-D 69.0 94.9 90.9 78.6 83.4
32-D 71.8 93.2 90.9 80.4 84.1
64-D 73.2 93.2 86.4 80.4 83.3
128-D 71.8 93.2 81.8 78.6 81.4
256-D 74.7 93.2 86.4 78.6 83.2
512-D 74.7 93.2 81.8 80.4 82.5
Table 7: Average accuracy with varying feature dimensions on LERF scenes
8 Evaluation on 3DOVS Dataset